OpenStudy (anonymous):

Write an equation for the sequence 5, 10, 17, 26

OpenStudy (zzr0ck3r):

?

OpenStudy (ivettef365):

you have 2^2 +1 = 5 3^2 + 1 = 10 4^2 + 1 = 17 5^2 + 1 = 26

OpenStudy (ivettef365):

so it is n^2+1